ON MEASURE AND HAUSDORFF DIMENSION OF JULIA

SETS FOR HOLOMORPHIC COLLET–ECKMANN MAPS

Feliks Przytycki

Abstract. Let f : C̄ → C̄ be a rational map on the Riemann sphere , such that for every f-critical point c ∈ J which forward trajectory does not contain any other critical point, |(f^{n})′(f(c))| grows exponentially fast (Collet–Eckmann condition), there are no parabolic periodic points, and else such that Julia set is not the whole sphere. Then smooth (Riemann) measure of the Julia set is 0.

For f satisfying additionally Masato Tsujii’s condition that the average distance of f^{n}(c) from the set of critical points is not too small, we prove that Hausdorff dimension of Julia set is less than 2. This is the case for f(z) = z^{2} + c with c real, 0 ∈ J, for a positive line measure set of parameters c.

Typeset by AMS-TEX